### 12 Dec 2023 v1

# Near-Optimal Differentially Private $k$-Core Decomposition

Abstract
Recent work by Dhulipala, Liu, Raskhodnikova, Shi, Shun, and Yu [${}^{+}$22] initiated the study of the $k$-core decomposition problem under differential privacy. They show that approximate $k$-core numbers can be output while guaranteeing differential privacy, while only incurring a multiplicative error of $(2+\eta)$ (for any constant $\eta>0$) and additive error of $\text{poly}(\log(n))/\varepsilon$. In this paper, we revisit this problem. Our main result is an $\varepsilon$-edge differentially private algorithm for $k$-core decomposition which outputs the core numbers with no multiplicative error and $O(\text{log}(n)/\varepsilon)$ additive error. This improves upon previous work by a factor of 2 in the multiplicative error, while giving near-optimal additive error.
With a little additional work, this implies improved algorithms for densest subgraph and low out-degree ordering under differential privacy. For low out-degree ordering, we give an $\varepsilon$-edge differentially private algorithm which outputs an implicit orientation such that the out-degree of each vertex is at most $d+O(\log{n}/{\varepsilon})$, where $d$ is the degeneracy of the graph. This improves upon the best known guarantees for the problem by a factor of $4$ and gives near-optimal additive error. For densest subgraph, we give an $\varepsilon$-edge differentially private algorithm outputting a subset of nodes that induces a subgraph of density at least ${D^{*}}/{2}-O(\text{log}(n)/\varepsilon)$, where $D^{*}$ is the density for the optimal subgraph.
1 Introduction
The $k$-core decomposition is a fundamental graph characteristic that offers deep insights into the structure of any graph. It is also an unique problem in the sense that it has managed to receive much more attention in recent years from the database and systems communities (for a comprehensive survey, see [MGPV20]) compared with the algorithms community (see [${}^{+}$22] and references therein). Such interests within the database and systems communities largely stem from the many applications of solutions to the problem to various practical tasks including community detection [AHDBV05, ELM18, GLM19], clustering and data mining [${}^{+}$11, SERF16], and various machine learning and graph analytic applications [KM17, DBS17, DBS18]. Due to the pace by which practical implementations of $k$-core decomposition is currently being developed in many applications involving sensitive user data, it is becoming ever more important to develop private algorithms for the problem.
The purpose of this paper is two-fold. First, we seek to fill in the gap for accurate, private and efficient algorithms for $k$-core decomposition. Second, we seek to contribute to the growing but nascent body of theoretical work on private $k$-core decomposition and related problems. The problem has such widespread applications because the $k$-core decomposition is a very efficient and accurate measure of the set of important, well-connected vertices in a graph. Such a set of vertices is colloquially known as the “core” of the graph. Formally, the $k$-core of a graph is the maximal induced subgraph of $G$ where each vertex in the induced subgraph has degree at least $k$. Since this definition captures a single core, the $k$-core decomposition problem (sometimes also called the coreness problem) is to identify the core number of every vertex $v$, which is defined to be the largest $k$ for which $v$ is in the $k$-core, but not in the $(k+1)$-core. [^1] We note that the $k$-core decomposition problem in the literature is sometimes used to refer to the problem of computing a hierarchy of $k$-cores based on graph connectivity [SP16]; note that it is not possible to emit such a hierarchy privately, so we (and prior work) focus on privately emitting core numbers for every vertex. Identifying the coreness values of every vertex is an important way of classifying the structural importance of each node in the graph.
Private $k$-core decomposition was introduced in the recent work of Dhulipala, Liu, Raskhodnikova, Shi, Shun, and Yu [${}^{+}$22] with differential privacy guarantees. Differential privacy [DMNS06] is the gold standard for privacy in modern data analytics. They study the problem in the very strong local edge model of differential privacy (LEDP). In particular, the local model guarantees the privacy of each individual adjacency list such that no edge is known by any additional party beyond the two vertices that are the endpoints of the edge. In other words, there is no notion of a trusted third-party in the local model of differential privacy. Although such a model provides strong privacy guarantees, it is often a restrictive model for a few reasons. First, it is inherently a distributed model where vertices send messages over several rounds. Thus, it is important that the number of rounds of communication is small (e.g. $\text{poly}(\log n)$). Another subtlety of the local model is that often the number of queries to a particular vertex needs to be limited in order to achieve good approximation guarantees; this limits the set of algorithms we may consider for the problem. By ensuring a small number of distributed rounds (and number of queries to any particular vertex), when translating such algorithms to the central model (run on a centralized server on one machine), the running time often becomes superlinear. In particular, the $k$-core decomposition algorithm of [${}^{+}$22], when directly translated to the central model, requires $\Omega(n^{2}\log n)$ time. Second, algorithms in the local model often have worse approximation guarantees than algorithms in the central model for the same problem. For the $k$-core decomposition problem, [${}^{+}$22] produces approximate core numbers for every vertex with multiplicative approximation error $2+\eta$ and $O(\text{poly}(\log n)/\varepsilon)$ additive error, for any constant $\eta>0$, in $O(\log n)$ rounds. However, as there is no lower bound showing that a $(2+\eta)$-multiplicative approximation is necessary for privacy, it was left open whether one can improve on this multiplicative approximation factor. Additionally, the additive approximation in [${}^{+}$22] requires a fairly large poly-logarithmic factor and it was also left open whether this result could be improved.
While the local model provides the strongest privacy guarantees, such strong privacy guarantees are sometimes unnecessary in cases where the graph is known by some external party anyways. (Take, for example, social network companies who already have access to the private network data.) In this paper, we revisit the differentially private $k$-core decomposition problem and provide improved utility guarantees in the central model of differential privacy. The central model assumes a trusted curator which has access to the private data and carefully processes it before releasing the privatized information. We work in the differential privacy models under edge-differential privacy. Our $k$-core decomposition algorithm in particular provides better utility guarantees than the recent work of [${}^{+}$22], providing an approximation with no multiplicative factor of the coreness values of every vertex with only an $O(\log n/\varepsilon)$ additive error with high probability. Our $k$-core decomposition algorithm is based on a novel formulation of the AboveThreshold procedure from the sparse vector technique which we call the multidimensional AboveThreshold technique (MAT). Like most existing algorithms for $k$-core, our algorithm is based on iterative peeling, i.e. iteratively removing the lowest degree vertices from the graph until all remaining vertices have large enough degree. The analysis of our algorithm leverages a new analysis of this iterative peeling process based on a multidimensional version of the sparse vector technique [${}^{+}$09].
1.1 Our Contributions
Given an undirected graph $G=(V,E)$, let $n=|V|$ be the number of vertices and $m=|E|$ be the number of edges in the graph. We can define the $k$-core of a graph and core numbers of the vertices as follows:
Definition 1.1.
A $k$-core of the graph is a maximal induced subgraph $H$ of $G$ where each vertex in $H$ has induced degree at least $k$. A vertex $v$ has core number (or coreness) $k$ if $v$ is in the $k$-core but not the $(k+1)$-core of $G$.
When outputting approximate core numbers of a graph, one usually only requires a multiplicative approximation. In our setting, the algorithms need to additionally incur additive error in order to preserve differential privacy. Our notion of approximate $k$-core decompositions is defined as follows:
Definition 1.2.
Given a vertex $v$, let $k(v)$ be the core number of $v$ and suppose the algorithm outputs $\hat{k}(v)$ as the approximate core number of $v$. Then we say that the algorithm gives a $(\phi,\zeta)$-approximation to the core numbers if we have $k(v)-\zeta\leq\hat{k}(v)\leq\phi\cdot k(v)+\zeta$ for all $v\in V$.
Our main result is a private algorithm for $k$-core decomposition which improves both the multiplicative and additive approximation guarantees of the algorithm in [${}^{+}$22]. We cleverly adapt a standard peeling algorithm for the $k$-core decomposition to satisfy $\varepsilon$-edge differential privacy and shows that the resulting algorithm achieves a $(1,O(\log{n}/\varepsilon))$-approximation to the core numbers. Since the maximum $k$-core number is within a factor of $2$ of the densest subgraph in any graph by the Nash-Williams theorem, the additive lower bound given by [NV21, FHS22] also translates to a $\Omega(\sqrt{\log(n)/\varepsilon})$ additive error lower bound for the $k$-core decomposition problem, so our result is nearly tight.
Theorem 1.3.
There is an algorithm which gives a $(1,O(\frac{\log(n)}{\varepsilon}))$-approximation to the core numbers that is $\varepsilon$-edge differentially private and achieves the approximation guarantees with high probability.
We also study the closely related densest subgraph problem, which we define and state our results for next. The density of a subgraph, and the densest subgraph problem are defined as follows:
Definition 1.4.
Given an undirected graph $G(V,E)$, and a subgraph $S=(V_{S},E_{S})$ of $G$ where $E_{S}$ is the induced subgraph of $G$ on $V_{S}$, the density of $S$ is defined to be $d(S)=|E_{S}|/|V_{S}|$. The densest subgraph problem is to find a subgraph $H_{\max}$ of $G$ which maximizes the density.
We note that we cannot output the exact densest subgraph without violating privacy; here it is known that $\Omega(\sqrt{\log(n)/\varepsilon})$ additive error is necessary [NV21, FHS22] for any $\varepsilon$-DP algorithm that succeeds with high probability in $n$. We define an approximate densest subgraph as follows:
Definition 1.5.
Let the optimal density of a densest subgraph in $G$ be $D^{*}$. A subgraph $H$ is a $(\phi,\zeta)$-approximation to the densest subgraph if $d(H)\geq\frac{D^{*}}{\phi}-\zeta$ for $\phi$, $\zeta\geq 1$.
Our main result for densest subgraph is given next, which follows from our $k$-core decomposition result.
Theorem 1.6.
There is an algorithm which gives a $(2,O(\log n/\varepsilon))$-approximation for the densest subgraph problem that is $\varepsilon$-edge differentially private and achieves the approximation guarantee with high probability.
We also give a differentially private algorithm for the low out-degree ordering problem (Definition 1.7). The low out-degree ordering problem seeks to output an ordering of the vertices in the graph to minimize the out-degree of any vertex when edges are oriented from vertices earlier in the ordering to later in the ordering. It is well-known that the best ordering(s) gives out-degree upper bounded by the degeneracy, $d$, of the graph.
Definition 1.7 ($(\phi,\zeta)$-Approximate Low Out-Degree Ordering).
Let $J=[v_{1},v_{2},\dots,v_{n}]$ be a total ordering of nodes in a graph $G=(V,E)$. The ordering $J$ is an $(\phi,\zeta)$-approximate low out-degree ordering if orienting edges from earlier nodes to later nodes in $J$ produces out-degree at most $\phi\cdot d+\zeta$.
Theorem 1.8.
There is an $\varepsilon$-edge differentially private algorithm that gives a $(1,O(\log n/\varepsilon))$-approximate low out-degree ordering of the vertices.
Furthermore, all of our algorithms can be implemented in near-linear time if we lose a multiplicative factor of $1+\eta$ for any given constant $\eta>0$. This nearly matches the best running times for sequential non-private $k$-core decomposition algorithms, while maintaining a near-optimal utility for private algorithms.
1.2 Related Work
The three problems our paper study, $k$-core decomposition, densest subgraph and low out-degree ordering, are closely related and have been studied side by side in the literature. For instance, the classic (non-private) $2$-approximation to the densest subgraph problem by Charikar [Cha00] works by performing an iterative greedy peeling process that is equivalent to computing the $k$-core decomposition, followed by post-processing. Subsequent work improved the approximation factor by using the multiplicative weight update framework [BGM14, SV20]. There has also been a large body of work on dynamic densest subgraph, with recent breakthrough results of Sawlani and Wang [SW20] and Chekuri and Quanrud [CQ22] yielding $O(\mathsf{poly}(\log n))$ time update while maintaining a $(1+\eta)$-approximate densest subgraph. The works of [${}^{+}$22, DKLV23] rely heavily on ideas from this literature.
On the private side, the first results for private densest subgraph began with the work of Nguyen and Vullikanti [NV21] and Farhadi et al. [FHO21]. [NV21] give algorithms which achieve a $O(2+\eta,\log(n)\log(1/\delta)/{\varepsilon})$-approximation under $(\varepsilon,\delta)$-differential privacy. On the other hand,  [FHO21] give algorithms which achieve a $(2+\eta,O({\log^{2.5}(n)\log(1/\sigma)}/{\varepsilon}))$-approximation with probability $1-\sigma$ again under $\varepsilon$-differential privacy. More recently, [${}^{+}$22] gave a $(1+\eta,O(\mathsf{poly}(\log n)/\varepsilon))$ approximation achieving $\varepsilon$-differential privacy by adapting the muliplicative weight update distributed algorithms of [BGM14, SV20] to the private setting; however, the private algorithm is not distributed and runs in $O((n+m)\log^{3}n)$ time. Both [NV21] and [FHO21] provide lower bounds on the additive error of any private densest subgraph algorithm; in particular, the lower bound of [FHO21] states that any $\varepsilon$-DP densest subgraph algorithm achieving high-probability accuracy guarantees must have additive error $\Omega(\sqrt{\log(n)/\varepsilon})$. Our densest subgraph result yields the same multiplicative factor as [NV21, FHO21] but improves the additive factor; on the other hand, we incur a worse multiplicative factor compared with [${}^{+}$22] but a better additive factor.
For $k$-core decomposition, to the best of our knowledge [${}^{+}$22] is the first work to study this problem; they give the first private algorithm achieving a $(2+\eta,O(\mathsf{poly}(\log n)/\varepsilon)$-approximation in the local edge-differential privacy setting by adapting a recently developed parallel locally-adjustable dynamic algorithm from [${}^{+}$22] to the private setting. To the best of our knowledge, their result is also the first $\varepsilon$-differential privacy result. Since their approach is based on a parallel algorithm, improving on this result, e.g., to achieve any $(2-\eta)$-approximation seems challenging due to a known $\mathsf{P}$-completeness result for degeneracy [AM84] (which reduces to $k$-core decomposition). We take a different approach in this paper, and adapt a variant of the classical peeling algorithm [MB83], which outputs exact core numbers with high probability, to the private setting. We note that this algorithm is not parallelizable, and indeed, the depth (i.e., longest chain of sequential dependencies) of this algorithm could be $\Omega(n)$ in the worst case.
Concurrent, Independent Work.
The recent concurrent, independent work of Dinitz et al. [DKLV23] provides a large set of novel results on $(\varepsilon,\delta)$-DP, $\varepsilon$-LEDP, and $(\varepsilon,\delta)$-LEDP densest subgraph, improving on the approximation guarantees of previous results. Their work is based on a recent result of Chekuri, Quanrud, and Torres [CQT22] and the parallel densest subgraph algorithm of Bahmani, Kumar, and Vassilvitskii [BKV12]. They offer improvements in the bounds in the central DP setting via the private selection mechanism of Liu and Talwar [LT19]. Although [DKLV23] also studies the densest subgraph problem (closely related to the $k$-core decomposition problem), their techniques are different from those presented in this paper; furthermore, any $c$-approximate algorithm for densest subgraphs translates to a $2c$-approximate algorithm for finding the degeneracy of the input graph (which is the value of the maximum $k$-core). Thus, no densest subgraph algorithm can obtain a better than $2$-approximation of the maximum $k$-core value.
2 Differential Privacy Background
We give some basic definitions and background on differential privacy in this section. We refer the interested reader to [${}^{+}$14] for proofs of these fundamental concepts. Since we will only be working with graphs, we will state these results in terms of edge differential privacy for graph algorithms; we will also use the terms differential privacy and edge differential privacy interchangeably. We start with the definition of edge-neighboring graphs and edge-differential privacy. Edge-neighboring graphs is a well-studied model for differential privacy (see e.g. [${}^{+}$23] for a survey of such results) and protects the privacy of connections with highly sensitive information like disease transmission graphs.
Definition 2.1.
We say that two graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ are edge-neighboring if $V_{1}=V_{2}$ and $|E_{1}\Delta E_{2}|=1$ (i.e., they have the same vertex set and they differ by exactly one edge). We often use the notation $G_{1}\sim G_{2}$ to denote that $G_{1},G_{2}$ are edge-neighboring.
Definition 2.2.
Let $\mathcal{G}$ denote the set of undirected graphs. We say an algorithm $\mathcal{M}:\mathcal{G}\to\mathcal{Y}$ is $(\varepsilon,\delta)$-edge differentially private if for all edge-neighboring graphs $G\sim G^{\prime}$ and every $S\subseteq\mathcal{Y}$, we have
$\Pr[\mathcal{M}(G)\in S]\leq\exp(\varepsilon)\cdot\Pr[\mathcal{M}(G^{\prime})\in S]+\delta.$
If $\delta=0$, we say $\mathcal{M}$ is $\varepsilon$-edge differentially private.
Next, we define two important properties of differential privacy: composition and post-processing. The composition property enables us to easily combine private subroutines into a larger algorithm which maintains some (weaker) privacy guarantees. The post-processing property guarantees that the output of a private algorithm cannot be post-processed to become less private without using information from the private data.
Lemma 2.3.
Let $\mathcal{M}_{1}:\mathcal{G}\to\mathcal{Y}_{1}$ and $\mathcal{M}_{2}:\mathcal{G}\to\mathcal{Y}_{2}$ be $\varepsilon_{1}$- and $\varepsilon_{2}$-edge differentially private mechanisms, respectively. Then $\mathcal{M}:\mathcal{G}\to\mathcal{Y}_{1}\times\mathcal{Y}_{2}$ defined by $\mathcal{M}=(\mathcal{M}_{1},\mathcal{M}_{2})$ is $(\varepsilon_{1}+\varepsilon_{2})$-edge differentially private.
Lemma 2.4.
Let $\mathcal{M}:\mathcal{G}\to\mathcal{Y}$ be an $\varepsilon$-edge differentially private mechanism. Let $f:\mathcal{Y}\to\mathcal{Z}$ be an arbitrary randomized mapping. Then $f\circ\mathcal{M}:\mathcal{G}\to\mathcal{Z}$ is still $\varepsilon$-edge differentially private.
3 Generalized Sparse Vector Technique
We introduce a multidimensional version of the AboveTreshold mechanism, which is the basis of the algorithms in the paper. In the standard AboveTreshold mechanism from Section 3.6 in [${}^{+}$14], we are given as input a threshold $T$ and an adaptive sequence of sensitivity $1$ queries $f_{1},f_{2},\ldots$, and the goal is to output the first query which exceeds the threshold. In the multidimensional version, we have a $d$-dimensional vector of thresholds $\vec{T}=(T_{1},\ldots,T_{d})$ and an adaptive sequence of sensitivity $\Delta$ vector-valued queries $\vec{f}_{1},\vec{f}_{2},\ldots$, and the goal is to output the first query for which the $j^{th}$ coordinate exceeds the threshold $T_{j}$, for each $j\in[d]$. We give a natural generalization of the 1-dimensional AboveThreshold mechanism in Algorithm 1 and give its privacy proof below.
Input: Private graph $G$, adaptive sequence of queries $\{\vec{f}_{i}\}$, threshold vector $\vec{T}$, privacy parameter $\varepsilon$, sensitivity $\Delta$.
Output: A sequence of responses $\{\vec{a}_{i}\}$ where $a_{i,j}$ indicates if $f_{i,j}(G)\geq\vec{T}_{j}$
1:  for $j=1,\ldots,d$ do
2:     $\hat{T}_{j}\leftarrow\vec{T}_{j}+\text{Lap}(2\Delta/\varepsilon)$
3:  end for
4:
5:  for each query $\vec{f}_{i}$ do
6:     for $j=1,\ldots,d$ do
7:        Let $\nu_{i,j}\leftarrow\text{Lap}(4\Delta/\varepsilon)$
8:        if $f_{i,j}(G)+\nu_{i,j}\geq\hat{T}_{j}$ then
9:           Output $a_{i,j}=\top$
10:           Stop answering queries for coordinate $j$
11:        else
12:           Output $a_{i,j}=\bot$
13:        end if
14:     end for
15:  end for
Algorithm 1 Multidimensional AboveThreshold (MAT)
Theorem 3.1.
Algorithm 1 is $\varepsilon$-differentially private.
Proof.
Fix $G\sim G^{\prime}$ to be arbitrary edge-neighboring graphs. We first define some notation for simplicity in the remainder of the proof. Let $\mathcal{A}(G)$ and $\mathcal{A}(G^{\prime})$ denote the random variable representing the output of Algorithm 1 when run on $G$ and $G^{\prime}$, respectively. The output of the algorithm is a realization of the random variables, where for each coordinate $j\in[d]$, the output is $a_{\cdot j}\in\{\top,\bot\}^{r(j)}$ and has the form that for all $i<r(j)$, $a_{i,j}=\bot$ and $a_{r(j),j}=\top$. Our goal will be to show that for any output $a$, we have
$\Pr[\mathcal{A}(G)=a]\leq\exp(\varepsilon)\cdot\Pr[\mathcal{A}(G^{\prime})=a].$
Observe that there are two sources of randomness in the algorithm: the noisy thresholds $\hat{T}_{j}$ for each $j\in[d]$ and the perturbations of the queries $\nu_{i,j}$ for each $j\in[d]$, $i\in[r(j)]$. We will fix the random variables $\nu_{i,j}$ for each $j\in[d]$ and $i<r(j)$ by simply conditioning on their randomness; for simplicity, we omit this from notation. The remaining random variables are $\hat{T}_{j}$ and $\nu_{r(j),j}$ for each $j\in[d]$; we will be taking probabilities over this randomness, but this will remain implicit in our notation. The main observation needed is that the output of $\mathcal{A}$ is uniquely determined by the iterations $r(j)$ of the first time we have $f_{i,j}(G)+\nu_{i,j}\geq\hat{T}_{j}$ for each $j\in[d]$. Thus, we can define the (deterministic) quantity
$g_{j}(G)\coloneqq\text{max}\{f_{i,j}(G)+\nu_{i,j}:i<r(j)\}$
for each $j\in[d]$ so that the event that $\mathcal{A}(G)=a$ is equivalent to the event where
$f_{r(j),j}(G)+\nu_{r(j),j}>\hat{T}_{j}\text{ and }\hat{T}_{j}\geq g_{j}(G)\text{ for each }j\in[d].$ (1)
Let $p(\cdot)$ and $q(\cdot)$ denote the density functions of the random vectors consisting of $\hat{T}(j)$ and $\nu_{r(j),j}$ for each $j\in[d]$. Given the above observation, we can write the following, where $\hat{T}_{j}\in(f_{r(j),j}(G)+\nu_{r(j),j},g_{j}(G))$ indicates the aforementioned event,
$\Pr[\mathcal{A}(G)=a]$ $=\Pr[\hat{T}_{j}\in(f_{r(j),j}(G)+\nu_{r(j),j},g_{j}(G))\text{ $\forall j\in[d]$}]$
[…]
Now, we make a change of variables. Define
$\vec{v}^{\prime}_{j}$ $=\vec{v}_{j}+g_{j}(G)-g_{j}(G^{\prime})+f_{r(j),j}(G^{\prime})-f_{r(j),j}(G)$
$\vec{t}^{\prime}_{j}$ $=\vec{t}_{j}+g_{j}(G)-g_{j}(G^{\prime})$
for each $j\in[d]$. Since we have conditioned on the randomness of $\nu_{i,j}$ for each $j\in[d]$ and $i<r(j)$, we have $\|\vec{v}^{\prime}-\vec{v}\|_{1}\leq 2\Delta$ and $\|\tilde{t}^{\prime}-\tilde{t}\|_{1}\leq\Delta$ since the vectors $f_{i\cdot}(G)$ have sensitivity $\Delta$, implying that the vectors $g_{\cdot}(G)$ also have sensitivity $\Delta$. Also observe that the change of variables only changes the vectors $\vec{v}$ and $\vec{t}$ by a constant (since we are still conditioning on the randomness of $\nu_{i,j}$) so $d\vec{v}\,d\vec{t}=d\vec{v}^{\prime}\,d\vec{t}^{\prime}$. Thus, we can apply the change of variables to rewrite (2) as:
[…]
Plugging in the definitions of $t^{\prime}_{j}$ and $v^{\prime}_{j}$ and simplifying, this is equivalent to
[…]
Now, recall that we showed $\|\vec{v}^{\prime}-\vec{v}\|_{1}\leq 2\Delta$ and $\|\vec{t}^{\prime}-\vec{t}\|_{1}\leq\Delta$. This implies that $p(\vec{t}^{\prime})\leq\exp(\varepsilon/2)\cdot p(\vec{t})$ and $q(\vec{v}^{\prime})\leq\exp(\varepsilon/2)\cdot q(\vec{v})$ by the properties of the Laplace distribution. We can thus upper bound the above as
[…]
By the definition of probability, the above is exactly equal to
$\exp(\varepsilon)\cdot\textstyle\Pr[\hat{T}_{j}\in(f_{r(j),j}(G^{\prime})+\nu_{r(j),j},g_{j}(G^{\prime}))\text{ $\forall j\in[d]$}].$
Finally, by our observation in (1), this is equal to
$\exp(\varepsilon)\cdot\Pr[\mathcal{A}(G^{\prime})=a].$
Combining the sequence of equalities and inequalities, we have shown that
$\Pr[\mathcal{A}(G)=a]\leq\exp(\varepsilon)\cdot\Pr[\mathcal{A}(G^{\prime})=a].$
Since $G,G^{\prime}$ were arbitrary edge-neighboring graphs, this completes the proof.
4 Differentially Private $k$-Core Decomposition
In this section, we give our improved algorithm for differentially private $k$-core decomposition. We first present a variant of the classical (non-private) algorithm for $k$-core decomposition in Section 4.1, and then show how to make it private in Section 4.2. Finally, we show how to implement it in nearly-linear time in Section 4.3, while incurring only $(1+\eta)$-multiplicative error.
4.1 A Variation of the Classical Algorithm
The classical peeling algorithm begins with the full vertex set $V$, which is the 00-core of the graph. Given the $(k-1)$-core, the algorithm computes the $k$-core via an iterative peeling process: the algorithm repeatedly removes all nodes $v$ for which the induced degree is less than $k$, and labels the nodes which remain as being part of the $k$-core. Running this for $k$ from 1 up to $n$ gives the full algorithm, the pseudocode of which is given in Algorithm 2.
Input: Graph $G=(V,E)$.
Output: $k$-core value of each node $v\in V$
1:  Initialize $V_{0}\leftarrow V$, $t\leftarrow 0$, $\hat{k}(v)\leftarrow 0$ for all $v\in V$
2:  for $k=1,\ldots,n$ do
3:     repeat
4:        $t\leftarrow t+1$, $V_{t}\leftarrow V_{t-1}$
5:        for $v\in V_{t-1}$ do
6:           if $d_{V_{t-1}}(v)<k$ then
7:              $V_{t}\leftarrow V_{t}-\{v\}$
8:           end if
9:        end for
10:     until $V_{t-1}-V_{t}=\emptyset$
11:
12:     Update the core numbers $\hat{k}(v)\leftarrow k$ for all nodes $v\in V_{t}$
13:  end for
Algorithm 2
Theorem 4.1.
For each $v\in V$, the output $\hat{k}(v)$ given by Algorithm 2 is the core number of $v$.
Proof.
We will inductively show that the algorithm recovers the $k$-core of the graph. The base case of $k=0$ is easy. Now, assume that the algorithm finds the true $(k-1)$-core. Let $V(k)$ denote the subset of nodes which aren’t removed in the iterative process for $k-1$ in Line 2. We have that each node $v\in V(k)$ has induced degree at least $k$ in $V(k)$, or else it would have been removed. Thus, we know that the core numbers $k(v)$ is at least $k$ for each $v\in V(k)$, so we have that $V(k)$ is a subset of the true $k$-core. Now, let $K$ denote the true $k$-core. Since the $k$-core is always a subset of the $(k-1)$-core by definition, each node $v\in K$ is in $V_{t}$ at the beginning of the iterations. Furthermore, we know that $v\in K$ is never removed from $V_{t}$ since the induced degree is always at least $k$ since $K\subseteq V_{t}$. Thus, we have that $V(k)$ is a superset of the true $k$-core as well. Thus, $V(k)$ is the true $k$-core for each $k$, so all nodes are labelled correctly. ∎
4.2 Private Implementation of the Algorithm
It is difficult to turn the classical algorithm into a differentially private one because it has $\Omega(n)$ iterations, which causes us to incur $\tilde{\Omega}(n)$ additive error when using basic composition (Lemma 2.3). In fact, this was cited in [${}^{+}$22] as the motivation for basing their algorithms on parallel/distributed algorithms for $k$-core decomposition, since those algorithms often have $\text{poly}\log({n})$ round-complexity. Our main observation is that the private version of Algorithm 2 can be analyzed as a special case of the Multidimensional AboveThreshold mechanism, so it doesn’t need to incur the $\Omega(n)$ additive error due to composition.
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$.
Output: An $(1,120\log(n)/\varepsilon)$-approximate $k$-core value of each node $v\in V$
1:  $V_{0}\leftarrow V$, $t\leftarrow 0$, $k=60\log{n}/\varepsilon$.
2:  Initialize $\hat{k}(v)\leftarrow 0$ for all $v\in V$
3:  for $v\in V$ do
4:     ~$\tilde{\ell}(v)\leftarrow\text{Lap}(4/\varepsilon)$
5:  end for
6:
7:  while $k\leq n$ do
8:     repeat
9:        $t\leftarrow t+1$, $V_{t}\leftarrow V_{t-1}$
10:        for $v\in V_{t-1}$ do
11:           $\nu_{t}(v)\leftarrow\text{Lap}(8/\varepsilon)$
12:           if $d_{V_{t-1}}(v)+\nu_{t}(v)\leq k+\tilde{\ell}(v)$ then
13:              $V_{t}\leftarrow V_{t}-\{v\}$
14:           end if
15:        end for
16:     until $V_{t-1}-V_{t}=\emptyset$
17:
18:     Update the core numbers $\hat{k}(v)\leftarrow k$ for all nodes $v\in V_{t}$
19:     $k\leftarrow k+60\log{n}/\varepsilon$
20:  end while
Algorithm 3
Theorem 4.2.
Algorithm 3 is $\varepsilon$-edge differentially private.
Proof.
We will show that Algorithm 3 is an instance of the multidimensional AboveThreshold mechanism, implying that it is $\varepsilon$-edge differentially private. Specifically, we will show that its output can be obtained by post-processing the output of an instance of Algorithm 1. Indeed, consider the instance of the multidimensional AboveThreshold mechanism where the input graph is $G$, the privacy parameter is $\varepsilon$, and the threshold vector $\vec{T}=\vec{0}$ is the zero vector. We will now inductively (and adaptively) define the queries.
For each iteration $t$, the $t^{th}$ query consists of the vector of $d_{V_{t-1}}(v)-k$ for each $v\in V$, where $V_{0}=V$ as in the algorithm; note that this matches with the queries on Line 12 of the algorithm. First, observe that the sensitivity of the vector containing $k-d_{V_{t-1}}(v)$ for each $v\in V$ is $\Delta=2$ for each iteration $t$, as needed in the algorithm. Next, observe that we can construct $V_{t}$ from $V_{t-1}$ using only the outputs $a_{t}(v)$ of the queries at the current iteration: given $v\in V_{t-1}$, we include $v$ in $V_{t}$ as well if and only if $a_{t}(v)=\bot$, which is equivalent to the condition in Line 12 of $d_{V_{t-1}}(v)+\nu_{t}(v)\leq k+\tilde{\ell}(v)$ [^2] Here, we implicitly assume that the randomness used by the algorithm and the AboveThreshold mechanism are the same. This is justified by conditioning on the random variables used by the two algorithms, proving differential privacy using the conditional probabilities. We then taking expectations over the conditional probabilities to conclude privacy for the full algorithm.. Thus, this is a feasible sequence of adaptive queries for the AboveThreshold mechanism.
Since the sequence of subsets $\{V_{t}\}$ are obtained by post-processing the outputs $a_{t}(v)$ of the AboveThreshold mechanism, they are $\varepsilon$-differentially private by Theorem 3.1 and Lemma 2.4. By applying post-processing again, the sequence of pairs $(V_{t},k_{t})$ for each iteration $t$ is also $\varepsilon$-differentially private since the sequence $k_{t}$ is public. Given the pairs $(V_{t},k_{t})$ for each iteration $t$, we can now recover the approximate core numbers which the algorithm outputs by setting the core numbers as $\hat{k}(v)=k_{t}$ for each node in $V_{t}-V_{t-1}$. It is easily verified that this gives the exact same output as Algorithm 3, so we have $\varepsilon$-differential privacy for the algorithm by applying post-processing (Lemma 2.4) once again. ∎
Theorem 4.3.
Algorithm 3 outputs $(1,\frac{120\log{n}}{\varepsilon})$-approximate core numbers $\hat{k}(v)$ with probability $1-O(\frac{1}{n^{2}})$.
Proof.
By the density function of the Laplace distribution, we know that we have $|\nu_{t}(u)|\leq\frac{40\log{n}}{\varepsilon}$ and |~$|\tilde{\ell}(u)|\leq\frac{20\log{n}}{\varepsilon}$ each with probability at least $1-\frac{1}{n^{5}}$. Since there are at most $O(n^{3})$ such random variables, taking a union bound over all nodes $u\in V$ and all iterations of the loops, we have the above guarantee with probability at least $1-{O}\left(\frac{1}{n^{2}}\right)$. We condition on the event that the above inequalities hold true for each $t\in[T]$ and each $u\in V$ for the remainder of the proof.
Fix an arbitrary iteration $k$ of the while loop starting on Line 7. Let $H$ be the set of nodes remaining in $V_{t}$ at the end of the while loop in Lines 8—16. We claim that ($i$) all nodes $u\in H$ have core number at least $k-\frac{60\log{n}}{\varepsilon}$ and ($ii$) all nodes $u\not\in H$ have core number at most $k+\frac{60\log{n}}{\varepsilon}$. To see ($i$), consider the subgraph $H$; we claim that each node $u\in H$ has induced degree at least $k-\frac{60\log{n}}{\varepsilon}$ in $H$. Indeed, since each node in $H$ was not removed in the final iteration of the while loop in Lines 8—16, we have that $d_{H_{t}}(u)+\nu_{t}(u)\geq k+\tilde{\ell}(u)$ for each $u\in H$. But since we have assumed $|\nu_{t}(u)|\leq\frac{40\log{n}}{\varepsilon}$ and |~$|\tilde{\ell}(u)|\leq\frac{20\log{n}}{\varepsilon}$, our desired bounds follow directly by the triangle inequality. To see ($ii$), let’s suppose for contradiction that the $k$-core value of $u$ is $k(u)>\ell+\frac{60\log{n}}{\varepsilon^{\prime}}$. Then there exists a subgraph $u\in K\subseteq V$ where the induced degree of each node $v\in K$ is $d_{K}(v)\geq k(u)$. But for such a subgraph $K$, the condition in Line 12 will always be true (again, because $|\nu_{t}(u)|\leq\frac{40\log{n}}{\varepsilon}$ and |~$|\tilde{\ell}(u)|\leq\frac{20\log{n}}{\varepsilon}$) so $K\subseteq H$. But since $u\in K$, this contradicts the fact that $u\not\in H$.
Using the above bounds, we can now prove our desired results. First, consider an arbitrary node $u\in V$ labeled $\hat{k}(u)$ within the while loop in Lines 7—20 and not relabeled later. Since $u$ was labeled $\hat{k}(u)$ in the current iteration, we have that $k(u)\geq\hat{k}(u)-\frac{60\log{n}}{\varepsilon^{\prime}}$. Since $u$ was not relabelled in the next iteration where the threshold was $k=\hat{k}(u)+\frac{60\log{n}}{\varepsilon}$, the node $u$ was removed from $V_{t}$ at that iteration. Consequently, we have $k(u)\leq k+\frac{60\log{n}}{\varepsilon}=\hat{k}(u)+\frac{120\log{n}}{\varepsilon}$ by what we just proved above. Since the output of our algorithm is $\hat{k}(u)$, the desired bounds in the theorem statement follow directly. Now, let’s consider an arbitrary node $u\in V$ labeled 00 at Line 2 at the beginning of the algorithm and not relabeled later. Since the node $u$ was not relabelled at the first iteration where $k=\frac{60\log{n}}{\varepsilon}$, it was removed from $V_{t}$ at that iteration, implying that $k(u)\leq\frac{120\log{n}}{\varepsilon}$. Hence, we have the desired approximation guarantees for all nodes. ∎
The above proof can be modified with appropriate constants to give the approximation with probability at least $1-\frac{1}{n^{c}}$ for any $c\geq 1$, thus yielding our desired with high probability result.
4.3 Efficient Implementation of the Algorithm
Though the current algorithm has strong utility guarantees, the running time is at least quadratic in $n$. We will make a slight modification to the algorithm, which will incur an additional multiplicative error of $1+\eta$ for an user chosen constant $\eta>0$, and show that a clever sampling trick can be used to implement the resulting algorithm in time nearly linear in the number of edges $m$. Via a similar proof as before, we can obtain an analogous set of privacy and utility guarantees.
Theorem 4.4.
Algorithm 4 has the following guarantees:
  • It is $\varepsilon$-edge differentially private.
  • It outputs $(1+\eta,O(\log{n}/\varepsilon))$-approximate core numbers with probability $1-O(\frac{1}{n^{2}})$.
Input: Graph $G=(V,E)$, multiplicative error $\eta>0$, privacy parameter $\varepsilon>0$.
Output: An $(1+\eta,60\log(n)/\varepsilon)$-approximate $k$-core value of each node $v\in V$
1:  $V_{0}\leftarrow V$, $t\leftarrow 0$, $k=60\log{n}/\varepsilon$.
2:  Initialize $\hat{k}(v)\leftarrow 0$ for all $v\in V$
3:  for $v\in V$ do
4:     ~$\tilde{\ell}(v)\leftarrow\text{Lap}(4/\varepsilon)$
5:  end for
6:
7:  while $k\leq n$ do
8:     repeat
9:        $t\leftarrow t+1$, $V_{t}\leftarrow V_{t-1}$
10:        for $v\in V_{t-1}$ do
11:           $\nu_{t}(v)\leftarrow\text{Lap}(8/\varepsilon)$
12:           if $d_{V_{t-1}}(v)+\nu_{t}(v)\leq k+\tilde{\ell}(v)$ then
13:              $V_{t}\leftarrow V_{t}-\{v\}$
14:           end if
15:        end for
16:     until $V_{t-1}-V_{t}=\emptyset$
17:
18:     Update the core numbers $\hat{k}(v)\leftarrow k$ for all nodes $v\in V_{t}$
19:     $k\leftarrow(1+\eta)\cdot k$
20:  end while
Algorithm 4
Now, we discuss the efficient sampling procedure. We claim that the naive implementation of the algorithm given in Algorithm 3 requires $\Omega(n^{2})$ runtime in the worst case. Let us focus on the do-while loop in Lines 8–16 of Algorithm 3. Clearly, the for-loop in Lines 10–15 iterates $\Theta(n)$ times for each iteration of the do-while loop. We claim that in the worst case, the do-while loop also requires $\Omega(n)$ iterations. Let us illustrate in the easy case where there is no noise (or equivalently, when we take $\varepsilon$ to infinity). Consider the path graph, and observe that the degree of each node is $2$ except the endpoints which have degree $1$. If the threshold $\ell$ is 2, then only the endpoints are removed and the resulting graph will again be a path graph, now of size $n-2$. Clearly, it will take $n/2$ iterations before the do-while loop terminates. This example can be generalized to the noisy version by increasing the separation of the degrees from $1$ up to $4\log{n}$.
We will give a more efficient way to implement the sampling so that the algorithm can run in $\tilde{O}(m+n)$ time. The idea is as follows: given a vertex $v$ and a fixed threshold $k+\tilde{\ell}$, we can calculate the probability $p$ that the vertex is removed at a given iteration of the do-while loop. In fact, this probability remains the same for vertex $v$ unless a neighbor of $v$ is removed from $H_{t}$. Thus, whenever a neighbor of $v$ is removed, we can sample a (geometric) random variable for when the vertex $v$ is peeled assuming no more neighbors of $v$ are removed. Whenever another neighbor of $v$ is removed, we have to resample this geometric random variable, but the number of times this resampling must be done over all vertices $v$ scales linearly with the number of edges $m$ rather than quadratically with the number of nodes $n$. This is detailed below. In Algorithm 5, $\text{Geom}(q)$ is the geometric distribution with parameter $q$.
    Input: Graph $G=(V,E)$, parameter $\varepsilon>0$, threshold $k+\tilde{\ell}(v)$ for each $v\in V$.
    Output: $V_{t}$ with same distribution as Lines 8–16 of Algorithm 3
1:  $t\leftarrow 0$, $V_{0}\leftarrow V$
[…]
4:  $\text{updated}\leftarrow V$
5:  repeat
6:     for $v\in\text{updated}$ do
[…]
8:        $q\leftarrow\Pr[d_{V_{t}}(v)+\text{Lap}(8/\varepsilon)\geq k+\tilde{\ell}(v)]$
[…]
11:     end for
12:
13:     $t\leftarrow t+1$, $V_{t}\leftarrow V_{t-1}$//this is only notational; we only store the set $H_{t}$ for a single $t$
14:
15:     for $v\in\text{to-remove}(t)$ do
16:        remove $v$ from $V_{t}$
17:        add all neighbors of $v$ in $V_{t}$ to updated//need to update remove-time if degree changes
18:     end for
19:  until $V_{t-1}-V_{t}=\emptyset$
Algorithm 5 An efficient implementation of Lines 8–16 in Algorithm 3
Lemma 4.5.
Algorithm 5 satisfies the following:
  • It gives the same distribution of $V_{t}$ as Lines 8–16 in Algorithm 3.
  • It can be implemented in $\tilde{O}(n+m)$ time.
Proof.
To see that the distribution is the same, we use induction on $t$. We claim that at each iteration $t$, the distribution of $V_{t}$ is the same for the two implementations of the algorithm. The statement is clear for $t=0$ since $V_{0}=V$ in both implementations. Now, assume that the distribution of $V_{t-1}$ is the same for both implementations; we will show that the distribution of $V_{t}$ conditioned on $V_{t-1}$ is also the same for both implementations. This will imply our desired result by induction.
Given $V_{t-1}$, let us compute that the probability that a node $v\in V_{t-1}$ is not included in $V_{t}$ in both implementations. Note that the event that a node is not included in $V_{t}$ is independent (conditioned on $V_{t}$) so showing that the probabilities are the same is sufficient to show the equivalence of (conditional) distributions. In the implementation in Algorithm 3, the probability is clearly just
$\Pr[d_{V_{t}}(v)+\text{Lap}(8/\varepsilon)\geq k+\tilde{\ell}(v)].$ (3)
Now, let’s consider the implementation in Algorithm 5. To understand when the node $v$ will be removed, we look at the final time $t^{\prime}$ that node $v$ is added to the set updated. The node $v$ will be removed at time $t$ if the remove-time(v) sampled by the algorithm is exactly $t$. The probability that this occurs conditioned on the event that $v$ was not removed at $V_{t-1}$ can be calculated as:
$\Pr[t^{\prime}+\text{Geom}(q)=t\mid t^{\prime}+\text{Geom}(q)\geq t-1]=\Pr[\text{Geom}(q)=1]=q,$
where the first equality holds by the memoryless property of the geometric distribution. Since no neighbors of $v$ were removed between iterations $t^{\prime}$ and $t$, the induced degrees $d_{V_{t^{\prime}}}(v)$ and $d_{V_{t}}(v)$ are the same. It is thus evident that the probability $q$ is the same as the expression in Eq. 3, proving the desired claim.
Now, we will prove that the implementation given in Algorithm 5 runs in near-linear time. First, we note that all sets are implemented as binary search trees so adding and removing elements requires only $O(\log{n})$ time. To analyze the time complexity, observe that the total run time of Lines 6–11 (in an amortized sense) scales with the total number of nodes added to the set updated (up to logarithmic factors). Since a node is only added to the set updated when one of its neighbors is removed, the total number of nodes added to updated during the entire algorithm scales with the number of edges. The time complexity of Lines 16 is $\tilde{O}(n)$ since each node is only removed once and the time complexity of Line 17 is $\tilde{O}(m)$ for the same reason as before. That completes the proof. ∎
5 Related Problems
We now discuss how our improved private $k$-core decomposition algorithm can be used to derive better private algorithms for densest subgraph and low out-degree ordering. We again note that all algorithms given here can be implemented in near-linear time using Lemma 4.5 while losing only $1+\eta$ in the multiplicative error; these details are omitted.
5.1 Differentially Private Densest Subgraph
We present our $(2,O(\log(n)/\varepsilon))$-approximate edge-differentially private densest subgraph algorithm in Algorithm 6. Our algorithm follows immediately from our $k$-core decomposition algorithm and the well-known folklore theorem that the density of the densest subgraph, $D(G)$, of graph $G$ falls between half of the maximum $k$-core value and the maximum $k$-core value. In fact, one can show that the $(k_{\max}/2)$-core (where $k_{\max}$ is the maximum value of a non-empty core) contains the densest subgraph in $G$ (see e.g. Corollary 2.2 of [SLDS23]). Furthermore, the $k_{\max}$-core gives a $2$-approximation of the densest subgraph.
We now use Algorithm 4 to obtain differentially private approximations of the $k$-core values of each vertex. Then, we find the maximum returned core value; we denote this value by $\hat{k}_{\max}$. Finally, we take the induced subgraph consisting of all vertices with approximate core values at least $\hat{k}_{\max}-\frac{c^{\prime}\log(n)}{\varepsilon}$ for sufficiently large constant $c^{\prime}$. This means that we include all $k_{\max}$ vertices, with high probability, and potentially some vertices with core values at least $k_{\max}-\frac{2c^{\prime}\log(n)}{\varepsilon}$. However, we prove that this still results in a good approximate densest subgraph in Theorem 5.1.
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$.
Output: An $(2,c^{\prime}\log(n)/\varepsilon)$-approximate densest subgraph $S\subseteq V$ for sufficiently large constant $c^{\prime}>0$
1:  Run Algorithm 3 with parameter $\varepsilon$ to obtain approximate core numbers $\hat{k}(v)$ for each $v\in V$
2:  Let $\hat{k}_{\max}\leftarrow\max_{v\in V}\left(\hat{k}(v)\right)$
3:  Find $S\leftarrow\{v\mid\hat{k}(v)\geq\hat{k}_{\max}-c^{\prime}\log(n)/\varepsilon\}$
4:  Return $S$
Algorithm 6
Theorem 5.1.
Algorithm 6 outputs an $\varepsilon$-edge differentially private $\left(2,O\left(\frac{\log(n)}{\varepsilon}\right)\right)$-approximate densest subgraph with probability $1-O\left(\frac{1}{n^{c}}\right)$ for any constant $c\geq 1$.
Proof.
The privacy of Algorithm 6 is guaranteed because we only do postprocessing on the differentially private output of Algorithm 3 and privacy is preserved under postprocessing (Lemma 2.4). Then, to prove the additive error, we first observe that with probability at least $1-O\left(\frac{1}{n^{c}}\right)$, the vertex set $S$ contains vertices with core numbers at least $k_{\max}-\frac{c^{\prime}\log(n)}{\varepsilon}$ for sufficiently large constant $c^{\prime}>0$. Then, we can calculate the minimum density of the graph induced by $S$ to be at least
$\frac{|S|\cdot(k_{\max}-2c^{\prime}\log(n)/\varepsilon)}{2|S|}=\frac{|S|\cdot k_{max}}{2|S|}-\frac{|S|\cdot 2c^{\prime}\log(n)}{2\varepsilon|S|}\geq\frac{k_{\max}}{2}-\frac{c^{\prime}\log(n)}{\varepsilon}.$
Since we know $k_{\max}/2\leq D(G)\leq k_{\max}$, it follows that $S$ is a $(2,O(\log(n)/\varepsilon))$-approximate densest subgraph. ∎
5.2 Differentially Private Low Out-Degree Ordering
[${}^{+}$22] defines the low out-degree ordering problem to be finding a differentially ordering where the out-degree of any vertex is minimized when edges are oriented from vertices earlier in the ordering to later in the ordering. Specifically, there exists an ordering of the vertices of any input graph where the out-degree is at most the degeneracy $d$ of the graph, where the degeneracy of the graph is equal to the maximum $k$-core value. The previous best differentially private algorithm for the problem gives out-degree at most $(4+\eta)d+O\left(\frac{\log^{3}n}{\varepsilon}\right)$. In this paper, we give a novel algorithm that obtains out-degree bounded by $d+O\left(\frac{\log n}{\varepsilon}\right)$, with high probability.
Input: Graph $G=(V,E)$, privacy parameter $\varepsilon>0$.
Output: An $\left(1,O\left(\log(n)/\varepsilon\right)\right)$-approximate $k$-core value of each node $v\in V$
1:  $V_{0}\leftarrow V$, $t\leftarrow 0$, $k=c^{\prime}\log{n}/\varepsilon$.
2:  Initialize $\hat{k}(v)\leftarrow 0$ for all $v\in V$
3:   Initialize ordered list $J$
4:  for $v\in V$ do
5:     ~$\tilde{\ell}(v)\leftarrow\text{Lap}(4/\varepsilon)$
6:  end for
7:
8:  while $k\leq n$ do
9:     repeat
10:        $t\leftarrow t+1$, $V_{t}\leftarrow V_{t-1}$
11:        for $v\in V_{t-1}$ do
12:           $\nu_{t}(v)\leftarrow\text{Lap}(8/\varepsilon)$
13:           if $d_{V_{t-1}}(v)+\nu_{t}(v)\leq k+\tilde{\ell}(v)$ then
14:              $V_{t}\leftarrow V_{t}-\{v\}$
15:               Append $v$ to the end of $J$
16:           end if
17:        end for
18:     until $V_{t-1}-V_{t}=\emptyset$
19:
20:     Update the core numbers $\hat{k}(v)\leftarrow k$ for all nodes $v\in V_{t}$
21:     $k\leftarrow k+c^{\prime}\log{n}/\varepsilon$
22:  end while
23:   Return $J$
Algorithm 7 Low Out-Degree Ordering
Our algorithm (given in Algorithm 7) is a simple modification of Algorithm 3 where each time we remove a vertex $v$ in Line 13 of Algorithm 3, we add $v$ to the end of our current ordering. The changes to Algorithm 3 are highlighted in blue. Thus, our ordering is precisely the order by which vertices are removed. We now prove that our algorithm for low out-degree ordering is $\varepsilon$-edge DP and gives an ordering for which the out-degree is upper bounded by $d+O\left(\frac{\log n}{\varepsilon}\right)$.
Theorem 5.2.
Our modified Algorithm 3 is $\varepsilon$-edge DP and gives a low out-degree ordering where the out-degree is upper bounded by $d+O\left(\frac{\log n}{\varepsilon}\right)$.
Proof.
We first prove that our modified algorithm is $\varepsilon$-edge DP. First, the proof of Theorem 4.3 ensures that the ratio of the probabilities that all vertices $v$ are removed at the same steps in edge-neighboring graphs $G$ and $G^{\prime}$ is upper bounded by $e^{\varepsilon}$. Thus, since the ordering of the vertices follows precisely the order by which vertices are removed, then, the ordering that is released is $\varepsilon$-edge DP.
Now, we prove the upper bound on the out-degree returned by the algorithm. Each removed vertex $v$ has induced degree, when removed, that is upper bounded by $k(v)+O\left(\frac{\log n}{\varepsilon}\right)$, with high probability, by Theorem 4.3. The induced degree when the vertex is removed is exactly equal to its out-degree. Then, since $d=\max_{v\in V}\left(k(v)\right)$, the out-degree of the returned ordering is at most $d+O\left(\frac{\log n}{\varepsilon}\right)$, with high probability. ∎
Acknowledgements
We thank anonymous reviewers for their comments and suggestions, and Jalaj Upadhyay for his feedback on the manuscript. George Z. Li and Laxman Dhulipala are supported by NSF award number CNS-2317194. Quanquan C. Liu is supported by an Apple Simons Research Fellowship.